Determine the pH of a 10-3 M HCl & 10-5 M HNO3 solution.
Determine the pH of 10-4 M NaOH & 10-2 M KOH solution.
Determine the pH of a solution with [H30+] of 6.7 E -4 mol/L & 2.5 E -2 M.
Find the [H30+] if the pH is 5; if the pH is 12.
The pH of an aqueous solution is 1.50; determine the [H30+] and the [OH-].
The pH of an aqueous solution is 3.67; determine the [H30+].
Determine the normal concentration of an unknown acid solution when 17.3 mL of this solution was used in titrating 31.2 mL of 0.125 N NaOH solution.
Determine the normality of a NaOH solution when 45.1 mL of a 0.100 N HCl solution is used in titrating 31.0 mL of a base.
If 20.0 mL of 0.0100 M HCl solution is required to neutralize 30 mL of NaOH; determine the molarity of the NaOH solution.
If 20.0 mL of 0.10 N NaOH solution is required to neutralize 12 mL of HCl; determine the normality of the HCl solution.
1
Expert's answer
2020-04-06T12:05:54-0400

1) HCl = H+ + Cl-, HNO3 = H+ + NO3-, pH = -lg[H+]

pH = 3 for 10-3 M HCl, pH = 5 for 10-5 M HNO3


2) pH = 14 - pOH, pOH = -lg[OH-], NaOH = Na+ + OH-, KOH = K+ + OH-

pH = 10 for 10-4 M NaOH, pH = 12 for 10-2 M KOH solution.


3) pH = 3.17 for 6.7 E -4 mol/L, pH = 1.60 for 2.5 E -2 M.


4) [H30+] = 10^(-5) if the pH is 5; [H30+] = 10^(-12) if the pH is 12.


5) [H30+] = 3.16*10^(-2) M, [OH-] = 3.16*10(-13) M for the pH of an aqueous solution 1.50.


6) [H30+] = 2.14*10^(-4) M for the pH of an aqueous solution 3.67.


7) C1*V1 = C2*V2, normal concentration of an unknown acid solution is 0.125*31.2/17.3 = 0.225 N.


8) normality of a NaOH solution is 45.1*0.100/31.0 = 0.132 N.


9) molarity of the NaOH solution is 20.0*0.0100/30 = 0.0067 M.


10) the normality of the HCl solution is 20*0.10/12 = 0.167 N.